OpenStudy (anonymous):

Write your own real-world story problem that requires adding or subtracting mixed numbers with unlike denominators.

OpenStudy (anonymous):

@ImNotGoodAtMath

OpenStudy (imnotgoodatmath):

Dylan added 2 3/4 gallons of water to her fish tank. She realized she needed more water, so she added 1 1/8 gallons more. How much water did she add all together? 2 3/4 + 1 1/8 = 11/4 + 9/8 = 22/8 + 9/8 = 31/8 = 3 7/8 gallons

OpenStudy (anonymous):

ty :D

OpenStudy (imnotgoodatmath):

:D

OpenStudy (anonymous):

also can you tell me is i got this question right

OpenStudy (imnotgoodatmath):

yeah

OpenStudy (anonymous):

Find the difference ten and two-ninths minus seven and two-thirds. Be sure to reduce to simplest form. four and five-ninths three and zero-sixths two and five-ninths three and four-ninths i think its D